Session: Feb 2003 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Andrea Orlati at Tue Feb 4 13:24:16 2003) Minor failures : FS
Prc file was not properly loaded so schedule couldn't initialize the system, in particular VC frequency was not correct. Problem solved starting from 9:31 UT. 2 scans were lost.

Onsala (Michael Lindqvist at Tue Feb 4 13:26:45 2003) Success :
No known problems. Tape shiped.

Metsahovi (Ari Mujunen at Tue Feb 4 14:12:09 2003) Minor failures : RECO HIGH PCAL
Constant dry snowfall. Recorder headstack 1 worn to -20um, increased vacuum from 5in to 9in, abandoning thick tape swapping. No phasecal injection. Tape shipped with DHL (448 2592 192) and recorded to TRACK. Putting logs on vlbeer shortly. Please use GPS data only starting from 16-Jan-2003, the date of last adjustment. (Current offset -32.5used, rate estimate is +5.26nsec/day or +0.061psec/s.)

Noto (P. Cassaro at Wed Feb 5 12:35:34 2003) Minor failures : POIN
A couple of times the antenna was slightly in late in reach the sources.

Jodrell_Bank (Zsolt Paragi (for A.G.) at Thu Feb 6 08:23:57 2003) Success :

Cambridge (Zsolt Paragi (for A.G.) at Thu Feb 6 08:27:03 2003) Minor failures : RECO
Due to a failure of the headstack, track 30 on the VLBA recorder has been remapped to track 34, i.e. third system track. Please take this into account when correlating. This will also be the case for the other Cm experiments. Also there was an error message from the Formatter suggesting an incorrect sampler assignment (see log).